OpenStudy (anonymous):

Which of the following points is a solution of 3x + 2y ≤ 8? A. (0, 5) B. (6, -2) C. (3, 0) D. (0, 3)

OpenStudy (mathmate):

Hint: Plug in (0,5) into 3x+2y=3(0)+2(5)=10 Since 10 is not \(\le\) 8, so A is not a valid solution. Try the same with the remaining choices to find the correct solution.